# Detecting Misuse of Security APIs: A Systematic Review

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2023conference: ACM Computing Survey; June, 2023; $c$s: Security and privacy  Software and application security11footnotetext: CREST: The Centre for Research on Engineering Software Technologies
@@ 1. Introduction
Security Application Programming Interfaces (APIs) are integral to modern software development that heavily serves billions of users through web or mobile apps. They offer developers specific functionalities, such as data encryption or access control, to address security concerns. Cryptography and SSL/TLS APIs, for example, are widely used to ensure data confidentiality and secure communications  (Georgiev et al., 2012). Although security APIs are highly beneficial for ensuring software security, their incorrect use, known as misuse, inadvertently leads to software vulnerabilities, posing a significant risk to overall system security and potentially exposing millions of users to sensitive data breaches and financial losses (Georgiev et al., 2012; Egele et al., 2013; Rahaman et al., 2019; Bianchi et al., 2018; Al Rahat et al., 2019; Krüger et al., 2019).
Figure 1 illustrates an SSL/TLS API misuse, where a developer uses the API to establish a secure connection with a server (Step ①) but chooses to trust all server hostnames (Step ②). By exploiting the misuse, a malicious actor can impersonate a valid server, intercept the communication between a user and the application, and obtain unauthorized access to user’s personal information (Step ③). Many developers do not fully understand the implications of trusting all hostnames while using an SSL/TLS API (Georgiev et al., 2012). The complex design of security APIs, poor documentation, and inadequate training significantly hamper developers’ understanding of the security APIs and contribute to the widespread misuse of these APIs. A comprehensive study of ten thousand Android applications by Krüger et al. (Krüger et al., 2019) reveals that nearly 95% of the applications contain at least one cryptography API misuse. Similarly, an analysis of over two thousand open-source Java projects on GitHub found that 72% of them suffer from at least one cryptography API misuse (Hazhirpasand et al., 2019).
@@ 2.2. Misuse of Security APIs
As digital transformation accelerates, the use of APIs, including security APIs, is growing. The increasing interconnectivity of systems and the proliferation of IoT devices expand the attack surface, making the correct use of security APIs more critical than ever. However, the research reveals a concerning trend of widespread misuse of these APIs among developers. (Krüger et al., 2019; Hazhirpasand et al., 2019), which can be attributed to several factors. Firstly, the complex and nuanced operation of security APIs coupled with the inherent complexity of API design can hinder developers’ understanding and correct implementation (Green and Smith, 2016). Security API use is further complicated by poorly written documentation, often lacking clear usage examples, and sometimes including insecure code examples. For instance, misuses were identified in code samples within the documentation of an OAuth API, which can mislead developers who copy them without understanding the security implication (Shernan et al., 2015). Additionally, developers often use unreliable forum posts for guidance, which can propagate instances of misuses (Fischer et al., 2017). Lastly, the lack of cybersecurity training, coupled with the evolving threat landscape surrounding security APIs, makes it challenging for developers to stay current on the latest best practices for security APIs (Assal and Chiasson, 2018).
@@ 3.2. Inclusion-Exclusion Criteria
We applied the inclusion and exclusion criteria outlined in Table 2 to filter the retrieved studies. We included studies on detecting misuse of security APIs in line with our research objectives and RQs (I1). The criteria were iteratively refined during the selection process to ensure an accurate selection of relevant papers. For instance, we excluded studies that focused on vulnerabilities in the internal design or implementation of security APIs (E1) or targeted detecting misuses of generic APIs (E2). We also introduced E3 and E4 as publications such as short papers lack sufficient information to address the RQs. Moreover, we adopted a venue assessment criterion (I2), similar to methods used in previous studies (Croft et al., 2022; Sabir et al., 2021; Sarker et al., 2024), to include only high-quality papers. This criterion involved including only papers published in ranked conferences according to the CORE ranking [^1] http://portal.core.edu.au/conf-ranks, and ranked journals according to the Scimago database [^2] https://www.scimagojr.com/journalrank.php. Both databases employ meticulous and comprehensive evaluation methodologies considering various factors to assess the quality and impact of venues. By utilizing these databases, we were able to identify high-quality papers effectively.
@@ 3.4. Data Extraction and Synthesis
Data Synthesis: We used descriptive statistics to analyze demographic information data items, while thematic analysis was used to analyze RQ-relevant data items. To conduct the thematic analysis, we followed the steps outlined in the guideline of study (Braun and Clarke, 2006). Firstly, we familiarized ourselves with the data by reading and examining the extracted data. Next, we generated initial codes to capture security APIs, misuses, detection techniques, and evaluation methods. Then, we searched for themes and generated potential themes for each data item by merging the corresponding initial codes based on their similarities. We reviewed the themes and mapped them iteratively to ensure all codes and themes were accurately allocated. Finally, we reviewed the synthesized results for each RQ and resolved any disagreements through regular meetings to finalize the answers to RQs.
@@ 3.5. Distribution of Studies
Figure 2.b depicts the distribution of the 69 primary studies included in this review across different years. No relevant studies were identified before 2012, but since then, there has been an upward trend in publications toward 2022. This increasing trend signifies a growing interest from the research community in security API misuse detection, highlighting the need for a systematic analysis of this evolving field. The figure also shows the number of papers based on the venue type including journals and conferences. Notably, a substantial majority—over 84%—of the selected papers (58 out of 69) were published in conferences. Figure 2.c provides further insight into publication venues, including journals and conferences with more than one publication. Security-specific venues are well-represented, with the ACM Conference on Computer and Communications Security (CCS) and the IEEE Symposium on Security and Privacy (S&P) emerging as the most popular venues with 9 and 5 publications, respectively. Software engineering venues are also present (e.g., Automated Software Engineering with 2 papers), demonstrating the interdisciplinary nature of this field.
@@ 4.1. API Taxonomy
width = rowsep = 0.2pt, colspec = Q[110]Q[120]Q[55]Q[307]Q[345], cell11-5 = c, cell2-131-3 = c, cell21 = r=5c, cell22 = r=5c, cell71 = r=2c, cell72 = r=2c, hline1, 2, 7,9, 10, 11, 12,13 = 1-5, hline3,4,5,6,8 = 3-5, APIs & Functionality Language Instances Study Refs
@@ 4.1.1. Cryptographic primitives APIs
Our survey covered 43 studies examining the usage of crypto APIs. Among these, 36 studies explored the usage of Java Cryptography Architecture (JCA), while Python PyCrypto API, C/C++ CommonCrypto API, JavaScript, and Go were each subject to 3, 3, 1, and 1 study respectively.
@@ 4.1.2. SSL/TLS APIs
Secure Sockets Layer (SSL) and Transport Layer Security (TLS) protocols are used to establish a secure channel for communication between a client and a server, protecting them from potential attacks like a Man-in-the-Middle (MitM). These protocols rely on cryptographic primitives to ensure authentication, confidentiality, and integrity of network messages. To establish and validate SSL connections, developers can utilize SSL/TLS APIs such as OpenSSL (ope, ), which encapsulate the details and functionalities of these protocols. A critical aspect of SSL connection establishment is authenticating the server. During the SSL handshake, shown in Figure 4.a, the server presents its public key certificate to the client as a means of authentication. It is essential that the client carefully verifies the authenticity of the server’s certificate to ensure the security of the SSL connection. The validation process involves certificate validation and hostname verification.
@@
In hostname verification, a client must verify that the hostname included in the certificate matches the hostname that the client is attempting to connect. This verification process prevents MitM attacks, where an attacker intercepts the communication between a client and server, and impersonates the server by sending false information to the client.
In our survey, 26 studies focused on the usage of SSL/TLS APIs, with 22 studies for Java APIs, e.g., Java Secured-Socket Extension (JSSE), and 6 studies for C/C++ APIs, e.g., OpenSSL.
Figure 4. (a) An overview of a simplified SSL handshake, (b) OAuth authorization code grant flow between user, relying party (RP) and service provider (SP), and (c) attestation process performed using Google SafetyNet Attestation API [S69]
@@ 4.1.3. OAuth APIs
Open Authorization (OAuth) (Hardt, 2012) is a popular authorization protocol that allows end-users to grant third-party websites or applications access to their private resources on a remote server without sharing their credentials. This process involves three major roles: the user or resource owner who owns protected resources, the Service Provider (SP) that hosts the resources, and Relying Party (RP), known as client application that uses the SP to obtain access to the user’s resources. Access tokens are issued by SPs to RPs with the owner’s approval for accessing protected resources. OAuth APIs provided by SPs such as Google, Twitter, or Facebook are used by developers to authenticate users or obtain access to users’ resources through their major accounts in SPs. Although OAuth was first introduced as an authorization framework, it has been widely adopted to implement Single-Sign-On authentication, making it difficult for developers to use it properly. To address this challenge, OpenID Connect (Sakimura et al., ) was introduced as an authentication framework based on OAuth. The OAuth specification (Hardt, 2012) defines four different protocol flows or grant types: (i) authorization code, (ii) implicit, (iii) resource owner password credentials, and (iv) client credentials. Figure 4.b depicts the process of authorization code grant which is the the most commonly used grant type. As shown in Figure 4.b, the process begins with a user sending a request to an RP to access a remote resource (step 1). The RP then redirects the user to the SP with an APP ID and an optional state parameter to bind this request (step 2). Next, the user authenticates with the SP and grants the RP’s requested permissions (step 3). The SP issues an authorization code and an optional state parameter to the user (step 4). The user is then redirected back to the RP’s redirection endpoint, where the request is rejected if the received state parameter mismatches the initial one (step 5). Next, the RP sends the authorization code and it’s secret (established during registration with the SP) to the SP to request an access token (step 6). The SP verifies the RP app by validating the App ID and app secret and then responds with an access token (steps 7). With the access token the RP requests user data from the SP, which is then shared with the RP accordingly (steps 8-9). In the implicit grant (which is simpler than the authorization code grant), in step 4, the SP directly responds with an access token instead of an authorization code, without authenticating the RP. Resource owner password credentials and client credentials grants are rarely used.
@@ 4.1.4. Fingerprint APIs
Our review identified one study [S37] that investigated the usage of Fingerprint API in Android apps. Both Google (Google, 2015) and OWASP (OWASP-MASTG, 2017) guidelines recommend using a fingerprint reader in conjunction with cryptographic operations for secure authentication. This involves using the fingerprint to unlock a cryptographic key protected by the Trusted Execution Environments (TEE), rather than just recognizing the user. TEE, an integral part of modern smartphones, can securely generate and store cryptographic keys. Combining TEE with fingerprint readers for Two-factor authentication provides strong security comparable to external hardware devices such as YubiKeys (YubiCo, 2017). To interact with the fingerprint sensor and verify whether a legitimate user has touched it, four essential steps are required to follow [S37]. That begins with generating a cryptographic key where developers specify key properties via parameters such as setting the user authentication required parameter to True ensuring key usability only after a legitimate user has touched the fingerprint reader. Next, the key is unlocked through user authentication. If a legitimate user touches the sensor, the cryptographic key is unlocked, triggering a series of callback functions. Developers can override the fingerprint callbacks to handle different scenarios based on user legitimacy. Once authenticated, the unlocked key can be used by an app to encrypt, decrypt, or sign data. Google recommends using a previously generated private key to sign a server-provided authentication token to authenticate, and then to send this token to the app’s remote backend (Google, 2015).
@@ 4.1.6. SafetyNet Attestation APIs
Attackers can alter an application’s behavior either by directly modifying the app or by obtaining root access to the host system and injecting malicious code. Hence, developers need to ensure their app’s code integrity and the client device’s status. Google offers the SafetyNet Attestation API (Google, 2020) to check the integrity of a device or application and detect compromised devices and tampered applications. Figure 4.c illustrates the attestation process using this API. The attest function, requiring a nonce and an API Key, triggers the attestation API. The nonce is generated by the application’s backend server and sent to the device upon attestation request (step 1). The API Key is created using Google’s API Console, a platform for developers to manage their Google APIs. When attestation is requested (step 2), Google Mobile Services (GMS) conducts several checks on the device, forwarding the results to Google’s server (step 3). Google’s server responds with signed attestation data (step 4), which GMS delivers to the client (step 5). The client app extracts a JSON Web Signature (JWS) from the data, sending it to the backend server for validation (step 6), followed by the client server verifying the JWS (step 7). In our review, one study [S69] analyzed the usage of attestation API in Android applications.
@@ 4.2.2. Variations in findings
As shown in Figure 6, there are wide ranges of misuse rates for some APIs. For instance, crypto misuse rates in Android applications range dramatically from 25% to 100%. Even code snippets from developer forums exhibit a wide range of misuse rates (48%-90%). Similar variations are observed for other APIs and platforms such as SSL misuse rates in Ubuntu, which vary significantly from 7% to 100%. There are various reasons for these substantial differences. We identified five key factors contributing to these variations including, dataset size, data source, temporal consideration, misuse scope, and detection techniques.
First, the size of the datasets analyzed in each study can vary considerably. Android studies, for example, encompass investigations targeting a mere 45 applications [S15] to massive analyses exceeding 500,000 apps [S36]. Second, the data for analysis has been collected from various sources. For instance, the study exploring developer forums, [S64], reported differing misuse rates—90% in OJC, 71% in GAD, and 48% in GASD, reflecting the diverse nature of these platforms. It is important to consider that GASD is not exclusively focused on programming, potentially contributing to the lower misuse rate observed. Third, studies were conducted across different years and utilized datasets collected at various time intervals. Security practices evolve over time, potentially affecting misuse prevalence. For example, a study [S12] compared datasets from 2012 and 2016 to understand changes in misuse prevalence. Their analysis revealed a decrease in specific misuse types, likely due to increased developer awareness of insecure practices over time. Conversely, new misuses might be discovered over time, leading to apparent increases in misuse rates. Fourth, studies also differ in their focus and scope of misuse. Some delve into a single, specific misuse, while others aim for a broader analysis. Additionally, the prevalence of different misuse types naturally varies. The specific types and prevalence of misuses identified will be discussed in detail in Section 5. Last but not least, the detection techniques employed in these studies can introduce variations in the reported misuse rates. Some techniques may have limitations, such as reporting false positives or non-exploitable misuses. For example, a study [S69] initially reported a 100% misuse rate of SafetyNet Attestation in Android apps integrating this API, but after conducting exploit attempts, found that only a small set of apps had exploitable misuses. Section 6 will dive deeper into the details of detection techniques.
Despite the varied findings, it is evident that security API misuse remains a significant challenge for developers and our study represents a timely effort to raise awareness about this critical aspect of secure software development and to provide fundamental knowledge for the development of effective detection approaches.

### v3

@@ Abstract.
copyright: acmcopyrightjournal: CSURjournalyear: 2025ccs: Security and privacy Software security engineering11footnotetext: CREST: The Centre for Research on Engineering Software Technologies
@@ 1. Introduction
Security Application Programming Interfaces (APIs) are integral to modern software development that serves billions of users through web or mobile apps. They offer developers specific functionalities, such as data encryption or access control, to address security concerns. Cryptography and SSL/TLS APIs, for example, are widely used to ensure data confidentiality and secure communications  (Georgiev et al., 2012). Although security APIs are highly beneficial for ensuring software security, their incorrect use, known as misuse, inadvertently leads to software vulnerabilities, posing a significant risk to overall system security and potentially exposing millions of users to sensitive data breaches and financial losses (Georgiev et al., 2012; Egele et al., 2013; Rahaman et al., 2019; Bianchi et al., 2018; Al Rahat et al., 2019; Krüger et al., 2019).
[…]
@@ 2.2. Misuse of Security APIs
As digital transformation accelerates, the use of APIs, including security APIs, is growing. The increasing interconnectivity of systems and the proliferation of IoT devices expand the attack surface, making the correct use of security APIs more critical than ever. However, the research reveals a concerning trend of widespread misuse of these APIs among developers. (Krüger et al., 2019; Hazhirpasand et al., 2019), which can be attributed to several factors. Firstly, the complex and nuanced operation of security APIs coupled with the inherent complexity of API design can hinder developers’ understanding and correct implementation (Green and Smith, 2016). Security API use is further complicated by poorly written documentation, often lacking clear usage examples, and sometimes including insecure code examples. For instance, misuses were identified in code samples within the documentation of an OAuth API, which can mislead developers who copy them without understanding security implications (Shernan et al., 2015). Additionally, developers often use unreliable forum posts for guidance, which can propagate instances of misuse (Fischer et al., 2017). Lastly, the lack of cybersecurity training, coupled with the evolving threat landscape surrounding security APIs, makes it challenging for developers to stay current on the latest best practices for security APIs (Assal and Chiasson, 2018).
@@ 3.2. Inclusion-Exclusion Criteria
We applied the inclusion and exclusion criteria outlined in Table 2 to filter the retrieved studies. We included studies on detecting misuse of security APIs in line with our research objectives and RQs (I1). The criteria were iteratively refined during the selection process to ensure an accurate selection of relevant papers. For instance, we excluded studies that focused on vulnerabilities in the internal design or implementation of security APIs (E1) or targeted at detecting misuses of generic APIs (E2). We also introduced E3 and E4 to exclude publications, such as short papers, that lack sufficient information to fully address the RQs. Moreover, we adopted a venue assessment criterion (I2), similar to methods used in previous studies (Croft et al., 2022; Sabir et al., 2021; Sarker et al., 2024), to include only high-quality papers. This criterion involved including only papers published in ranked conferences according to the CORE ranking [^1] http://portal.core.edu.au/conf-ranks, and ranked journals according to the Scimago database [^2] https://www.scimagojr.com/journalrank.php. Both databases employ meticulous and comprehensive evaluation methodologies, considering various factors to assess the quality and impact of venues. By utilizing these databases, we were able to identify high-quality papers effectively.
@@ 3.4. Data Extraction and Synthesis
Data Synthesis: We used descriptive statistics to analyze demographic information data items, while thematic analysis was used to analyze RQ-relevant data items. To conduct thematic analysis, we followed the steps outlined in the guideline by (Braun and Clarke, 2006). Firstly, we familiarized ourselves with the data by reading and examining the extracted data. Next, we generated initial codes to capture security APIs, misuses, detection techniques, and evaluation methods. Then, we searched for themes and generated potential themes for each data item by merging the corresponding initial codes based on their similarities. We reviewed the themes and mapped them iteratively to ensure all codes and themes were accurately allocated. Finally, we reviewed the synthesized results for each RQ and resolved any disagreements through regular meetings to finalize the answers to RQs.
@@ 3.5. Distribution of Studies
Figure 2.b depicts the distribution of the 69 primary studies included in this review across different years. No relevant studies were identified before 2012, but since then, there has been an upward trend in publications toward 2022. This increasing trend signifies a growing interest from the research community in security API misuse detection, highlighting the need for a systematic analysis of this evolving field. The figure also shows the number of papers based on the venue type, including journals and conferences. Notably, a substantial majority—over 84%—of the selected papers (58 out of 69) were published in conferences. Figure 2.c provides further insight into publication venues, including journals and conferences with more than one publication. Security-specific venues are well-represented, with the ACM Conference on Computer and Communications Security (CCS) and the IEEE Symposium on Security and Privacy (S&P) emerging as the most popular venues with 9 and 5 publications, respectively. Software engineering venues are also present (e.g., Automated Software Engineering with 2 papers), demonstrating the interdisciplinary nature of this field.
@@ 4.1. API Taxonomy
width = rowsep = 0.25pt, colspec = Q[110]Q[120]Q[55]Q[307]Q[345], cell11-5 = c, cell2-131-3 = c, cell21 = r=5c, cell22 = r=5c, cell71 = r=2c, cell72 = r=2c, hline1, 2, 7,9, 10, 11, 12,13 = 1-5, hline3,4,5,6,8 = 3-5, APIs & Functionality Language Instances Study Refs
@@ 4.1.1. Cryptographic primitives APIs
Our survey covered 43 studies examining the usage of crypto APIs. Of these, 36 focused on the Java Cryptography Architecture (JCA), while the Python PyCrypto API and the C/C++ CommonCrypto API were each examined in 3 studies, and JavaScript and Go were each covered in 1 study.
@@ 4.1.2. SSL/TLS APIs
Secure Sockets Layer (SSL) and Transport Layer Security (TLS) protocols are used to establish a secure channel for communication between a client and a server, protecting them from potential attacks like a Man-in-the-Middle (MitM). These protocols rely on cryptographic primitives to ensure the authentication, confidentiality, and integrity of network messages. To establish and validate SSL connections, developers can utilize SSL/TLS APIs such as OpenSSL (ope, ), which encapsulate the details and functionalities of these protocols. A critical aspect of SSL connection establishment is authenticating the server. During the SSL handshake, shown in Figure 4.a, the server presents its public key certificate to the client as a means of authentication. It is essential that the client carefully verifies the authenticity of the server’s certificate to ensure the security of the SSL connection. The validation process involves certificate validation and hostname verification.
@@
In hostname verification, a client must verify that the hostname included in the certificate matches the hostname that the client is attempting to connect to. This verification process prevents MitM attacks, where an attacker intercepts the communication between a client and server, and impersonates the server by sending false information to the client.
In our survey, 26 studies focused on the usage of SSL/TLS APIs, with 22 studies for Java APIs, e.g., Java Secure Socket Extension (JSSE), and 6 studies for C/C++ APIs, e.g., OpenSSL.
Figure 4. (a) An overview of a simplified SSL handshake, (b) OAuth authorization code grant flow between user, relying party (RP), and service provider (SP), and (c) attestation process performed using Google SafetyNet Attestation API [S69]
@@ 4.1.3. OAuth APIs
Open Authorization (OAuth) (Hardt, 2012) is a popular authorization protocol that allows end-users to grant third-party websites or applications access to their private resources on a remote server without sharing their credentials. This process involves three major roles: the user or resource owner who owns protected resources, the Service Provider (SP) that hosts the resources, and Relying Party (RP), known as client application that uses the SP to obtain access to the user’s resources. Access tokens are issued by SPs to RPs with the owner’s approval for accessing protected resources. OAuth APIs provided by SPs such as Google, Twitter, or Facebook are used by developers to authenticate users or obtain access to users’ resources through their major accounts in SPs. Although OAuth was first introduced as an authorization framework, it has been widely adopted to implement Single-Sign-On authentication, making it difficult for developers to use it properly. To address this challenge, OpenID Connect (Sakimura et al., ) was introduced as an authentication framework based on OAuth. The OAuth specification (Hardt, 2012) defines four different protocol flows or grant types: (i) authorization code, (ii) implicit, (iii) resource owner password credentials, and (iv) client credentials. Figure 4.b depicts the process of authorization code grant, which is the most commonly used grant type. As shown in Figure 4.b, the process begins with a user sending a request to an RP to access a remote resource (step 1). The RP then redirects the user to the SP with an APP ID and an optional state parameter to bind this request (step 2). Next, the user authenticates with the SP and grants the RP’s requested permissions (step 3). The SP issues an authorization code and an optional state parameter to the user (step 4). The user is then redirected back to the RP’s redirection endpoint, where the request is rejected if the received state parameter mismatches the initial one (step 5). Next, the RP sends the authorization code and its secret (established during registration with the SP) to the SP to request an access token (step 6). The SP verifies the RP app by validating the App ID and app secret and then responds with an access token (steps 7). With the access token, the RP requests user data from the SP, which is then shared with the RP accordingly (steps 8-9). In the implicit grant (which is simpler than the authorization code grant), in step 4, the SP directly responds with an access token instead of an authorization code, without authenticating the RP. Resource owner password credentials and client credentials grants are rarely used.
@@ 4.1.4. Fingerprint APIs
Our review identified one study [S37] that investigated the usage of Fingerprint API in Android apps. Both Google (Google, 2015) and OWASP (OWASP-MASTG, 2017) guidelines recommend using a fingerprint reader in conjunction with cryptographic operations for secure authentication. This involves using the fingerprint to unlock a cryptographic key protected by the Trusted Execution Environments (TEE), rather than just recognizing the user. TEE, an integral part of modern smartphones, can securely generate and store cryptographic keys. Combining TEE with fingerprint readers for Two-factor authentication provides strong security comparable to external hardware devices such as YubiKeys (YubiCo, 2017). To interact with the fingerprint sensor and verify whether a legitimate user has touched it, four essential steps are required to follow [S37]. That begins with generating a cryptographic key where developers specify key properties via parameters such as setting the user authentication required parameter to True, ensuring key usability only after a legitimate user has touched the fingerprint reader. Next, the key is unlocked through user authentication. If a legitimate user touches the sensor, the cryptographic key is unlocked, triggering a series of callback functions. Developers can override the fingerprint callbacks to handle different scenarios based on user legitimacy. Once authenticated, the unlocked key can be used by an app to encrypt, decrypt, or sign data. Google recommends using a previously generated private key to sign a server-provided authentication token to authenticate, and then to send this token to the app’s remote backend (Google, 2015).
@@ 4.1.6. SafetyNet Attestation APIs
Attackers can alter an application’s behavior either by directly modifying the app or by obtaining root access to the host system and injecting malicious code. Hence, developers need to ensure their app’s code integrity and the client device’s status. Google offers the SafetyNet Attestation API (Google, 2020) to check the integrity of a device or application and detect compromised devices and tampered applications. Figure 4.c illustrates the attestation process using this API. The attest function, requiring a nonce and an API Key, triggers the attestation API. The nonce is generated by the application’s backend server and sent to the device upon attestation request (step 1). The API Key is created using Google’s API Console, a platform for developers to manage their Google APIs. When attestation is requested (step 2), Google Mobile Services (GMS) conducts several checks on the device, forwarding the results to Google’s server (step 3). Google’s server responds with signed attestation data (step 4), which GMS delivers to the client (step 5). The client app extracts a JSON Web Signature (JWS) from the data, sending it to the backend server for validation (step 6), followed by the client server verifying the JWS (step 7). In our review, one study [S69] analyzed the usage of the attestation API in Android applications.
@@ 4.2.2. Variations in findings
As shown in Figure 6, there are wide ranges of misuse rates for some APIs. For instance, crypto misuse rates in Android applications range dramatically from 25% to 100%. Even code snippets from developer forums exhibit a wide range of misuse rates (48%-90%). Similar variations are observed for other APIs and platforms, such as SSL misuse rates in Ubuntu, which vary significantly from 7% to 100%. There are various reasons for these substantial differences. We identified five key factors contributing to these variations, including dataset size, data source, temporal consideration, misuse scope, and detection techniques.
First, the size of the datasets analyzed in each study can vary considerably. Android studies, for example, encompass investigations targeting a mere 45 applications [S15] to massive analyses exceeding 500,000 apps [S36]. Second, the data for analysis has been collected from various sources. For instance, the study exploring developer forums, [S64], reported differing misuse rates—90% in OJC, 71% in GAD, and 48% in GASD, reflecting the diverse nature of these platforms. It is important to consider that GASD is not exclusively focused on programming, potentially contributing to the lower misuse rate observed. Third, studies were conducted across different years and utilized datasets collected at various time intervals. Security practices evolve over time, potentially affecting misuse prevalence. For example, a study [S12] compared datasets from 2012 and 2016 to understand changes in misuse prevalence. Their analysis revealed a decrease in specific misuse types, likely due to increased developer awareness of insecure practices over time. Conversely, new misuses might be discovered over time, leading to apparent increases in misuse rates. Fourth, studies also differ in their focus and scope of misuse. Some delve into a single, specific misuse, while others aim for a broader analysis. Additionally, the prevalence of different misuse types naturally varies. The specific types and prevalence of misuses identified will be discussed in detail in Section 5. Last but not least, the detection techniques employed in these studies can introduce variations in the reported misuse rates. Some techniques may have limitations, such as reporting false positives or non-exploitable misuses. For example, a study [S69] initially reported a 100% misuse rate of SafetyNet Attestation in Android apps integrating this API, but after conducting exploit attempts, it found that only a small set of apps had exploitable misuses. Section 6 will dive deeper into the details of detection techniques.
Despite the varied findings, it is evident that security API misuse remains a significant challenge for developers, and our study represents a timely effort to raise awareness about this critical aspect of secure software development and to provide fundamental knowledge for the development of effective detection approaches.